OpenStudy (anonymous):

Ivanna has scored 85, 81, 82, and 94 on her previous four tests. What score does she need on her next test so that her average (mean) is 84 ?

OpenStudy (anonymous):

Averages are found by adding up all the pieces and dividing by the number of terms added...

OpenStudy (anonymous):

... so (85 + 81+ 82 + 94 + x)/5 = 84 and solve

OpenStudy (anonymous):

...solve by multiplying both sides by 5 and then subtracting all the numbers on the left from each side

OpenStudy (anonymous):

which is....?

OpenStudy (anonymous):

wait why multiply

OpenStudy (anonymous):

because 5 is in the denominator (dividing) on the left, so you "undo" that by multiplying. ...there's another method...

OpenStudy (anonymous):

whats the other method

OpenStudy (anonymous):

ok so i came out with 1710=420????/ what did i do wrong

OpenStudy (anonymous):

oh wait is the answer 68?

OpenStudy (anonymous):

NEVER mind i got the right answer thanks!

OpenStudy (anonymous):

Sorry, I got pulled away, the other method is to keep track of how far above or below 84 each number is: 85 - 84 = 1 81 - 84 = -3 82 - 84 = -2 94 - 84 = 10 Then add up all of these, being careful about negatives = 1 - 3 - 2 + 10 = 6 So Ivanna has 6 extra points, which means she can get 84 - 6 = 78